The decomposition of phosphine is a first order process with a half life of 37.9 s. What is the rate constant for this reaction?

0.0264 s-1
0.0381 s-1
0.0183 s-1
54.7 s-1
26.3 s-1
Answer #1

Use the concentration time reaction and for first order reactions half life is given by

t1/2 = 0.693/k for first order process

k= 0.693/37.9s

k=0.0183 s^-1